Find the Cartesian equation of the locus of the point P representing the complex
number z. Sketch the locus of P in each case.

(a) 2|z + 1| = |z − 2|

(b) |(z + i )/ z - 5 - 2i | = 1

(c) Im (z + (9/z ) ) = 0
1
Expert's answer
2020-05-27T18:48:38-0400

(a) "2|z + 1| = |z \u2212 2|"

Solution:

Let "z=x+iy"

Substitute z in (a)

"2|x+iy+1|=|x+iy-2|"

"2\sqrt{\smash[b]{x+iy+1}}=\sqrt{\smash[b]{x+iy-2}}"

"2\sqrt{\smash[b]{x+1+iy}}=\sqrt{\smash[b]{x-2+iy}}"

"2\sqrt{\smash[b]{(x+1)^2+y^2}}=\sqrt{\smash[b]{(x-2)^2+y^2}}"

"2((x+1)^2+y^2)=(x-2)^2+y^2"

"2x^2+4x+2+2y^2=x^2-4x+4+y^2"

"2x^2-x^2+2y^2-y^2+4x+4x=4-2"

Answer: "x^2+y^2+8x=2"




(b) "|(z + i )\/ z - 5 - 2i | = 1"

Solution:

Cross multiply equation to get;

"|z+i|=|z-5-2i|"

Let "z=x+iy"

Substitute for z;

"x+iy+i=x+iy-5-2i"

"x+i(y+1)=x-5+i(y-2)"

"\sqrt{\smash[b]{x^2+(y+1)^2}}=\sqrt{\smash[b]{(x-5)^2+(y-2)^2}}"

"x^2+y^2+2y+1=x^2-10x+25+y^2-4y+4"

Simplify;

"x^2-x^2+y^2-y^2+10x+2y+4y=25+4-1"

Answer : "10x+6y=28"




(c) "Im (z + \frac{9}{z}) = 0"

Solution:

Let "z =x+iy"

Substitute for z in the expression;

"x+iy+\frac{9}{x+iy}"

Simplify; "\frac{x+iy(x+iy)}{x+iy}+\frac{9}{x+iy}"

"\frac{(x+iy)(x+iy)+9}{x+iy}"

"\frac{x^2+2ixy-y^2+9}{x+iy}"

Rationalize the denominator;

"\frac{x^2+2ixy-y^2+9}{x+iy}*\frac{x-iy}{x-iy}"

"\frac{x^2(x-iy)+2ixy(x-iy)-y^2(x-iy)+9(x-iy)}{x^2+y^2}"

"\frac{x^3-ix^2y+2ix^2y+2xy^2-y^2x+iy^3+9x-9iy}{x^2+y^2}"

Put the like terms together;

"\frac{x^3+2xy^2-y^2x+9x-ix^2y+2ix^2y+iy^3-9iy}{x^2+y^2}"

"\frac{x^3+2xy^2-y^2x+9x}{x^2+y^2} +\frac{i(x^2y+y^3-9y)}{x^2+y^2}"

"Im" part states;

"\frac{(x^2y+y^3-9y)}{x^2+y^2}=0"

"x^2y+y^3-9y=0"

"y(x^2+y^2-9)=0"

"x^2+y^2-9=0"

Answer: "x^2+y^2=9"